### v2

# Two-dimensional electronic spectroscopy in the condensed phase using equivariant transformer accelerated molecular dynamics simulations

Abstract
Two-dimensional electronic spectroscopy (2DES) provides rich information about how the electronic states of molecules, proteins, and solid-state materials interact with each other and their surrounding environment. Atomistic molecular dynamics simulations offer an appealing route to uncover how nuclear motions mediate electronic energy relaxation and their manifestation in electronic spectroscopies, but are computationally expensive. Here we show that, by using an equivariant transformer-based machine learning architecture trained with only  2500 ground state and  100 excited state electronic structure calculations, one can construct accurate machine-learned potential energy surfaces for both the ground-state electronic surface and excited-state energy gap. We demonstrate the utility of this approach for simulating the dynamics of Nile blue in ethanol, where we experimentally validate and decompose the simulated 2DES to establish the nuclear motions of the chromophore and the solvent that couple to the excited state, connecting the spectroscopic signals to their molecular origin.
Extracting information about the interplay between electronic and nuclear dynamics in complex condensed phase materials, ranging from light-harvesting complexes to quantum dots, is often greatly enhanced by going beyond linear electronic spectroscopy. Two-dimensional electronic spectroscopyHybl et al. (1998); Hybl, Ferro, and Jonas (2001) (2DES) is a nonlinear spectroscopic technique that provides access to features that encode population dynamics within and between states as well as the couplings of these states distributed across a range of excitation and emission frequencies. Characterizing these features and attributing them to specific molecular motions, particularly in condensed phase environments where these can be highly influenced by the structure and dynamics of the environment (e.g. solvent), is challenging because of the complexity of these systems. Understanding these observations has been deeply enriched by leveraging atomistic simulation. It is therefore important to develop efficient simulation methods that can accurately predict 2DES signals capturing phenomena such as anharmonic and vibronic effects, thermal broadening, and electronic interactions with the environment. Capturing these effects requires generating atomistic molecular dynamics (MD) on timescales of 50 to 100 picoseconds, along with the corresponding electronic excitation energies, to sample the nuclear fluctuations of both the chromophore and its environment. Assuming an ab initio treatment of both the dynamics and the environment, achieving this requires hundreds of thousands of ground and excited state electronic structure calculations for systems that may exceed 500 atoms owing to the desire to include potential charge transfer and polarization between the chromophore and its condensed phase environment. Such a large number of calculations is typically intractable at anything but the lowest levels of electronic structure theory, such as semiempirical methods.
Recent work has suggested opportunities to use machine learning (ML) approaches to provide a tractable route to efficiently compute and understand electronic properties ranging from electronic excitation energies in the gas phase to condensed phase linear and 2D electronic spectra. For gas-phase systems these ML approaches have been used to predict orbital energy gapsFuchs et al. (2020); Luo et al. (2023); Liao and Smidt (2023); Cignoni et al. (2024) and excitation energiesRamakrishnan et al. (2015); Pronobis et al. (2018); Ghosh et al. (2019); Mazouin, Schöpfer, and von Lilienfeld (2022); Entwistle et al. (2023); Baker, Pasini, and Hauck (2024) for databases of molecules at their optimized geometries. To obtain gas phase linear spectra that incorporate the shape of the electronic absorption peaks arising from molecular geometries beyond their optimized structure, recent work has trained ML models to predict electronic energy gaps, i.e. the difference in energy between the electronic states, at multiple geometries and combined these with the ensemble approachMukamel (1995); Crespo-Otero and Barbatti (2012); Isborn et al. (2012); Ge et al. (2015); Marenich, Cramer, and Truhlar (2015); Milanese et al. (2017); Zuehlsdorff et al. (2017) to compute linear spectra Ye et al. (2019); Xue, Barbatti, and Dral (2020); Lu et al. (2020); Westermayr and Marquetand (2020); Zou et al. (2023); Vinod et al. (2023); Petrusevich et al. (2023). ML potential energy surfaces have also been trained to the quality required to perform nonadiabatic dynamics simulations to study electronic relaxation through conical intersections in gas phase systems Chen et al. (2018); Westermayr et al. (2019); Axelrod, Shakhnovich, and Gómez-Bombarelli (2022); Shakiba and Akimov (2024). Machine learned potentials also show considerable promise in condensed phase studies such as in the construction of model Hamiltonians for large networks of coupled chromophores Häse et al. (2016); Farahvash et al. (2020); Cignoni, Cupellini, and Mennucci (2023); Betti et al. (2024), calculating ensemble approach linear electronic absorption spectra by machine learning electronic energy gaps in implicit solvent Bononi et al. (2020); Chen et al. (2022), and using electrostatic machine learning embedding to accurately capture environmental and vibronic effectsZinovjev and Curutchet (2025).
Our recent work has sought to employ ML models to compute the excited state electronic energy gaps for chromophores in condensed phase environments enabling the combination of a dynamical treatment of linear and 2D electronic spectra via the second-order cumulant approachMukamel (1985, 1995); Zuehlsdorff et al. (2019), allowing for the inclusion of vibronic and anharmonic effects Zuehlsdorff et al. (2019), combined with an explicit treatment of the solventChen et al. (2020). This work has allowed us to show that one can train an accurate ML model of the electronic energy gap using only $\sim$2000 energies Chen et al. (2020) or just $\sim$300 electronic energy gap gradients Chen et al. (2023a). By exploiting transfer learning we have also shown that, starting from ML energy gap surfaces trained using computationally affordable excited-state electronic structure methods including time-dependent density-functional theory (TDDFT) or configuration interaction singles (CIS), one can adapt the model to the more accurate but more expensive equation-of-motion coupled cluster with singles and doubles (EOM-CCSD) level of theory using only 100s of energiesChen et al. (2023a). This has allowed us to show that the explicit treatment of the hydrogen bonds (H-bonds) to the anionic green fluorescent protein chromophore in liquid water combined with a high-level wavefunction-based description of the excited state electronic structure is essential to capture the experimentally observed breadth of the linear electronic absorption spectrum and that this, in turn, leads to large changes in the 2DES signals Chen et al. (2023a).
Here we show that using an Equivariant Transformer (EqT) architecture Vaswani et al. (2017); Thölke and de Fabritiis (2021); Pelaez et al. (2024), which naturally encodes permutational, rotational, and translational symmetries within the machine-learned model, one can capture both the ground state electronic potential energy and electronic energy gap surfaces to simulate atomistic MD, the excitation energy gaps, and the corresponding linear and 2D electronic absorption spectra of a solvated chromophore using the energies and gradients obtained from only 2546 ground state and 90 excited state electronic structure calculations (SI Sec. IB-ID) compared to $\sim$100,000 of both without the use of ML. We demonstrate this approach by simulating the electronic spectroscopy of Nile blue in ethanol and comparing it to recent 2DES experimentsSon, Mosquera-Vázquez, and Schlau-Cohen (2017). In doing so, we reveal how atomistic MD simulations can be used to unravel the rich information present in 2DES signals and highlight the additional understanding that one can extract about the role of the atomistic environment. The EqT approach offers significant improvements over our previous work as an EqT model trained with less than 100 excited state calculations can achieve a root mean squared error (RMSE) of less than 10 meV compared to 1000’s of training calculations achieving RMSEs of 50-80 meV using the atom-centered neural network architecture based on Chebyshev polynomial descriptorsArtrith, Urban, and Ceder (2017) that we previously employedChen et al. (2020). The EqT architecture thus offers an extremely data-efficient route to fully atomistic, condensed-phase predictions of linear and 2DES spectra. Avoiding the tour-de-force ab initio molecular dynamics and excited state electronic structure calculations previously required will enable tighter feedback loops between experiment and simulation (SI Sec. IE). Upon establishing the efficiency of this approach, we then compare our simulations of the linear, 2DES, and pump-probe spectra to recent experiments of Nile blue in ethanol to unravel how solvation both drives electronic energy relaxation and enhances coupling to specific chromophore vibrational motions.
Figure 1: Convergence of the spectra predicted by the EqT models of the electronic excitation energy gap. Top: Simulated linear electronic absorption spectrum of Nile blue in ethanol as a function of the number of configurations (configs) used to train the EqT electronic energy gap model. The intensity maxima of all the spectra are shifted to be at 0 cm-1 to allow for easier comparison of convergence of the line shapes. The 90 and 180 configuration models produce graphically indistinguishable line shapes. Bottom: Simulated 2DES spectra with experimental pulse spectral profile applied (SI Sec. II) at a time delay of $t_{2}=600~{\rm fs}$ as a function of the number of training configurations used for the EqT energy gap model. The zero frequency of the x and y-axes are defined in the same way as the linear spectra above. Each of the three models give very similar spectra with only a minimal increase in breadth noticeable from 45 to 90 configurations.
To calculate linear and 2DES spectra, we use the second-order cumulant approach, Mukamel (1985, 1995); Zuehlsdorff et al. (2019) which requires obtaining a one-time correlation function of electronic energy gaps from an MD trajectory. To train an ML surface on which to perform the ground state dynamics, we began with a 50 ps density-functional tight-binding (DFTB)Porezag et al. (1994); Seifert, Porezag, and Frauenheim (1996) trajectory of Nile blue solvated by 175 ethanol molecules, selected configurations, calculated their energies and gradients with DFT using the revPBE exchange-correlation functional, trained an EqT model, and then iteratively added more configurations until the machine-learned potential was stable and achieving a gradient RMSE error on the test set of 44 meV/Å (SI Sec. IB). To calculate the ground to excited state energy difference, we sampled configurations from the ground-state EqT dynamics, calculated CAM-B3LYP/6-31G* excitation energies and excited state gradients, treating the Nile blue and, on average, 118 ethanol molecules at the same level of theory using TeraChem v1.9 Seritan et al. (2021) (SI Sec. IC), and used these values to train a second EqT model. We used the final ground state EqT model to generate 1 ns of dynamics then used the excitation energy EqT model to predict the energy gaps for every configuration in that trajectory, yielding a trajectory of 500,000 energy gap predictions trained using fewer than 200 excited state electronic structure calculations (SI Sec. ID).
Fig. 1 shows the convergence of the simulated linear absorption and 2DES spectra for Nile blue in ethanol as the number of configurations used to train the EqT electronic energy gap are increased. From this, one can see that for the linear absorption spectra (top), even when only 45 configurations are used, the shape of the spectrum is close to convergence while the results obtained from 90 and 180 training configurations are graphically indistinguishable. The RMSE on the test set falls from 30 meV to 24 meV to 12 meV as the number of training points is increased across this range, suggesting that errors below $\sim$20 meV in the energy gap manifest negligibly in the linear spectrum. This is not surprising given that the width of the absorption spectrum spans 0.6 eV, which is 50 times larger than the RMSE in the EqT model with 180 training configurations. The bottom panel of Fig. 1 shows the convergence of the 2DES as the number of configurations on which the electronic energy gap model was trained is increased. Similar to the linear absorption spectrum, rapid convergence is observed with the model trained on 45 configurations, giving a slightly narrower 2D spectrum and those with 90 and 180 configurations being indistinguishable. The EqT approach therefore allows one to use $\sim$100 excited state electronic structure calculations to build an electronic energy gap model that yields converged linear and 2D electronic spectra.
Given the converged spectra obtained from the EqT model, we now compare these spectra with recent experiments. The top panel of Fig. 2 shows the simulations and experimentally obtained linear absorption spectrum of Nile blue in ethanolSon, Mosquera-Vázquez, and Schlau-Cohen (2017). To allow easier comparison of the shape of the spectrum, the simulated spectrum has been shifted such that its absorption maximum is aligned with the experiment. The energy gap model was trained using TDDFT with the CAM-B3LYP exchange-correlation functional and 6-31G basis set, which is known to systematically overestimate electronic excitation gaps of organic chromophoresLaurent and Jacquemin (2013); Shao et al. (2019) by 100s of meV, and in this case overestimates it by 0.32 eV (the spectral maximum is 1.97 eV experimentally and 2.29 eV from simulation). One could improve the description of the position of the absorption maximum by using transfer learning to higher level wavefunction methodsChen et al. (2023a, b, c), but since our primary focus here is the mechanisms by which solvent interactions give rise to the spectral shape, we apply a uniform 0.32 eV shift of the linear and 2DES spectra to allow for easier comparison with experiment. With the shift applied, our simulated linear absorption spectrum yields good agreement with the experiment (Fig. 2) with a small underestimate of the width and high-frequency tail of the spectrum. The slightly narrower spectrum may be due to TDDFT with CAM-B3LYP underestimating the effect on the excitation energies upon H-bonding with the solvent, which has recently been demonstrated for the green fluorescent protein chromophore in water where EOM-CCSD was shown to correctly capture this effectChen et al. (2023a), or it may be due to non-Condon fluctuations, which have also recently been shown to produce additional broadeningWiethorn et al. (2023).
To better understand the origins of the shape of the linear spectrum, we can analyze the spectral density, $J(\omega)$, which can be obtained fromMukamel (1985, 1995); Zuehlsdorff et al. (2019) the time correlation of the energy gap fluctuations $\delta U$ relative to the mean electronic energy gap,
$J(\omega)=\frac{\beta\omega}{2}\int_{-\infty}^{\infty}dt~e^{i\omega t}\langle\delta U(t)\delta U(0)\rangle.$ (1)
[…]
Figure 2: Comparison of simulated linear spectrum with experiment and spectral density with SERS. Top: Experimentally obtained (gray, SI Sec. IA) and simulated (blue) linear absorption spectra of the Nile blue chromophore in ethanol. Shown in the inset is the chemical structure of the Nile blue cation. Experiment and simulation are separately shifted (by 1.97 and 2.29 eV) so their peak maxima match at 0 $\mathrm{cm}^{-1}$ to allow comparison of the spectral line shapes. Bottom: Simulated spectral density ($J(\omega)$, blue) indicates which nuclear motions couple to the electronic excitation and experimentally obtained SERS signalAuguié et al. (2012) (gray) as a comparison. The SERS spectrum is shifted vertically so all signal is non-negative then normalized to a maximum value of one.
2DES spectra present a more strenuous test of simulation than linear absorption spectra owing to the richer details they encode. Fig. 3 compares the 2DES spectra obtained from simulation to that from experiment. The simulated spectra are shifted by the same amount used for the linear spectra and we have multiplied the signal by the experimental pulse spectral profile to account for the window of frequencies observable in the experiment (SI Sec. II). The features and general shape of the 2DES obtained from the experiment and simulation are in good agreement, except for the negative feature observed in the experimental signal to the right of the main peak. Such negative features can arise from excited state absorption (ESA) processes but transient absorption experiments suggest that the negative feature observed here is likely an artifact from the phasing process. Lewis et al. (2010); Son, Mosquera-Vázquez, and Schlau-Cohen (2017); Dostál et al. (2018); Mewes et al. (2021). The simulated signal does not have any negative intensity because our simulation only includes the ground and first excited states, so there are no higher electronic states to absorb into. While the negative features observed at higher excitation frequencies in the experimental 2DES likely arise from phasing artifacts, the experimental pump-probe spectra (SI Sec. IV) show negative intensity that arises from ESA above $\sim 2000~\mathrm{cm}^{-1}$. These pronounced negative intensity contributions from ESA should appear as negative intensity on the 2DES emission axis since the integral of the excitation axis onto it yields the time-dependent pump-probe spectra.Mukamel (1995); Jonas (2003) However, these contributions lie above 2000 $\mathrm{cm}^{-1}$ and hence they are not observed as they lie outside the range of observable frequencies due to the experiment’s pulse spectral profile. Despite this, ESA still affects the experimental 2DES by shifting the position of the peak maximum to a lower frequency than that of the linear absorption spectrum. This shift in the 2DES maximum is not captured in the simulations that do not include ESA.
Figure 3: Comparison between experimental and simulated 2DES spectra. Top: 2DES obtained from experimentSon, Mosquera-Vázquez, and Schlau-Cohen (2017) at delay times of 81, 307, and 600 fs. Both excitation and emission frequencies are shifted by 1.97 eV as in Fig. 2. Bottom: The simulated spectra at the same delays. Both axes are shifted by 2.29 eV and have the experimental pulse spectral profile applied (SI Sec. II).
Having established the accuracy of the simulated spectra, we can now use them to decompose the 2DES signal into its different contributions to uncover the role of solvent coupling to the excited state. The top panels of Fig. 4 show the decomposition of the simulated signal into the contributions arising from the ground state bleach (GSB) and stimulated emission (SE) processes in which the chromophore population evolves on its ground or excited electronic state, respectively, during the $t_{2}$ delay. These processes are challenging to distinguish experimentally because both have the same wave vectors but are trivial to separate in the simulations since they arise from different terms in the response functionBiswas et al. (2022).
Figure 4: Obtaining the Stokes shift and dynamic Stokes shift from our 2DES simulations. Top panels: The simulated 2DES spectra with the experimental pulse spectral profile applied is decomposed into its SE and GSB components at two time delays (left column: $t_{2}=0$ fs and right column: $t_{2}=1800$ fs). These components are identical at $t_{2}=0$ fs and then separate as the delay time is increased due to the stabilization of the excited state. The weighted means (blue and purple markers) are obtained by averaging the data between the black vertical lines and calculating the weighted mean emission frequency of the resulting distribution. Triangles and circles are used to show the position at $t_{2}=0~\mathrm{fs}$ and $t_{2}=1800~\mathrm{fs}$ respectively. Bottom panel: The position of the weighted means of the 2DES spectra for the SE and GSB components as a function of the delay time. The separation of the two lines shown at each time is used to calculate the dynamic Stokes shift and the long time value (shown with double-headed arrow), 1226 $\mathrm{cm}^{-1}$, is the Stokes shift.
As the system undergoes longer $t_{2}$ delays, the SE peak shifts to lower energies since the chromophore and solvent rearrange to stabilize the excited electronic state. In contrast, the GSB peak arises from evolution on the ground state and shifts by $\sim$450 $\mathrm{cm}^{-1}$ to higher energies in the first 50 fs (SI Sec. V). By tracking the weighted mean position of the emission frequency of the spectra corresponding to the GSB and SE processes (purple and blue lines respectively in Fig. 4), the dynamics of the excited state stabilization by the motions of the chromophore and solvent can be examined. The bottom panel of Fig. 4 shows the time evolution of the positions of the GSB and SE. As the $t_{2}$ time delay increases, the SE signal decreases in energy, leading to a larger separation between it and the GSB signal. The relaxation of the SE position can be decomposed into fast ($\sim$50 fs) and slow ($\sim$1 ps) components with the former corresponding to intramolecular relaxation within the existing solvation environment while the latter component corresponds to broader conformational rearrangements and relaxation of the solvation environment around Nile blue. The GSB and SE position of the weighted means oscillate with a $\sim$55 fs period corresponding to the prominent 590 $\mathrm{cm}^{-1}$ ring breathing mode in the spectral density. The gap shown in Fig. 4, $\nu(t_{2})$, can be used to calculate the dynamic Stokes shift,Yan and Mukamel (1990); Mukamel (1995) $S(t_{2})$, according to
$S(t_{2})=\frac{\nu(t_{2})-\nu(\infty)}{\nu(0)-\nu(\infty)}.$ (2)
Within linear response, the dynamic Stokes shift is proportional to the energy gap time correlation function, with the corresponding Fourier transform used to construct the spectral densityMukamel (1995).
The difference between the GSB and SE in the long-time limit $\nu(t_{2}=1800~{\rm fs})$ in Fig. 4 (marked with an arrow between the circles) can be measured as a shift of 1226 $\mathrm{cm}^{-1}$. This shift is an underestimate since we applied the experimental pulse spectral profile to our simulated result, which narrows the range of emission frequencies that contribute significantly to the weighted mean. In particular, the SE peak extends to lower frequencies than the laser used in the experiment can fully capture. When simulated spectra are used without the experimental pulse spectral profile applied (SI Sec. V), the predicted shift increases to 1524 $\mathrm{cm}^{-1}$. We note that peak maxima are often used to measure the splitting between the GSB and SELee et al. (2017). However, due to the large asymmetry in the peaks, this approach leads to a $\sim$2-fold reduction of the shift to 529 $\mathrm{cm}^{-1}$ with the pulse spectral profile applied and 623 $\mathrm{cm}^{-1}$ without.
The Stokes shift, the difference between the energy at which a molecule absorbs and fluoresces light, is defined using the maxima of the peaks. Experimentally, we measured this to be 1010 $\mathrm{cm}^{-1}$ (SI Sec. VI), consistent with previous values in the literatureBaumann et al. (2001); Jose and Burgess (2006); Ghanadzadeh Gilani et al. (2012) ranging from 891 to 956 $\mathrm{cm}^{-1}$. However, due to the asymmetries in the absorption and florescence spectra, which arise from phenomena such as vibronic effects, the weighted mean frequency provides a better way to characterize the position of the distribution. Using the weighted mean of the experimental spectra gives a value of 1870 $\mathrm{cm}^{-1}$ (SI Sec. VI), which is significantly larger than the Stokes shift obtained from the spectral maxima. The difference in the positions between the weighted mean absorption and fluorescence spectra can be obtained from the simulations by reweighting the thermally sampled ground state configurations to account for their probabilities on the excited state (SI Sec. VII).Zwanzig, Rauh, and Thorn (1954); Tuckerman (2010); Ceriotti et al. (2012) Evaluating this from our simulations gives a shift of 1776 $\mathrm{cm}^{-1}$, in good agreement with the experimental value obtained using the same approach (1870 $\mathrm{cm}^{-1}$).
The reorganization energy, $\lambda$, which characterizes the strength of the coupling between the electronic states and nuclear motions, is given by,
$\lambda=\frac{1}{\pi\hbar}\int_{0}^{\infty}\mathrm{d}\omega~\frac{J(\omega)}{\omega}$ (3)
where $J(\omega)$ is the spectral density. Based on our simulations, the reorganization energy is 924 $\mathrm{cm}^{-1}$. Under the second-order cumulant approximation, the reorganization energy is half the Stokes shift in the limit of Gaussian absorption and emission line shapes and equals half the difference between the average of the excitation energy in thermal equilibrium as measured on the ground- and excited-state potential energy surface, respectively.Mukamel (1995) From our simulations, using Eq. 3 the reorganization energy is 924 $\mathrm{cm}^{-1}$ and thus the Stokes shift is 1824 $\mathrm{cm}^{-1}$, which is consistent with the value of 1776 $\mathrm{cm}^{-1}$ obtained from the simulations by reweighting and the experimentally obtained difference between the weighted means of the absorption and fluorescence spectra of 1870 $\mathrm{cm}^{-1}$.
To isolate the microscopic origin of the stabilization of the excited state, we now consider the effect of removing the electronic contribution of the solvent from the simulated predictions and its effect on the reorganization energyLu et al. (2021); Myers et al. (2024). To do this, we took the training configurations used previously, which were generated from the ground state dynamics of Nile blue in ethanol, and removed (stripped) all the solvent from those configurations. We then ran excited state electronic structure calculations on these solvent-stripped configurations, which were then used to train an EqT electronic energy gap model. This model was then used to compute the electronic energy gaps for the whole Nile blue in ethanol trajectory, considering only the chromophore configurations. The solvent-stripped spectral density (Fig. 5) was then computed from these energy gaps. The resulting spectral density thus retains the same nuclear motions of the chromophore and, therefore, dynamic and structural changes that arise from it being in solution, but does not include the electronic effects of the solvent on modulating the electronic energy gap between the ground and excited state.
Figure 5: Analysis of the changes to the reorganization energy and spectral density upon stripping solvent. Top: Difference between the ground and S1 excited state electron density of Nile blue shown with two ethanol molecules hydrogen bonded to the primary amine (-NH2) group. The difference density reveals significant changes to the electron density across the entire Nile blue molecule. Middle: $\Lambda(\omega)$ for the solvated (blue) and solvent-stripped (orange) simulations along with the difference between them (green). The values at $\omega=2000~\mathrm{cm}^{-1}$ for the solvated and solvent-stripped simulations (890 $\mathrm{cm}^{-1}$ and 592 $\mathrm{cm}^{-1}$ respectively) are their reorganization energies. These reorganization energies have significant contributions from low frequency motion below 100 $\mathrm{cm}^{-1}$ (brown shaded), a sharp increase just below 600 $\mathrm{cm}^{-1}$, and in the 1100-1550 $\mathrm{cm}^{-1}$ region (shaded gray). However, the difference arises primarily from the two shaded regions only. Bottom: The spectral densities of the solvated (blue) and solvent-stripped (orange) simulations. The solvated spectral density is normalized such that the maximum value is 1 and the solvent-stripped spectral density is normalized with the same factor. The derivative of the difference in $\Lambda(\omega)$, $\frac{d}{d\omega}\Delta\Lambda(\omega)$ is also plotted (dashed green) to highlight the largest contributions to the deviation between solvated and solvent-stripped simulations.
The changes upon stripping the solvent are shown in Fig. 5 where we computed the following quantity,
$\Lambda(\omega)=\frac{1}{\pi\hbar}\int_{0}^{\omega}\mathrm{d}\omega^{\prime}~\frac{J(\omega^{\prime})}{\omega^{\prime}}.$ (4)
[…]
$\frac{d\Delta\Lambda(\omega)}{d\omega}=\frac{J_{solvated}(\omega)-J_{stripped}(\omega)}{\omega}$ (5)
where $\Delta\Lambda(\omega)=\Lambda_{solvated}(\omega)-\Lambda_{stripped}(\omega)$, $\Lambda_{solvated}(\omega)$ and $\Lambda_{stripped}(\omega)$ denote the $\Lambda(\omega)$ from the solvated and solvent-stripped simulations respectively. $J_{solvated}(\omega)$ and $J_{stripped}(\omega)$ denote the corresponding spectral densities. From this one can see that the intense ring breathing mode of the central heteroatom-containing ring (587 $\mathrm{cm}^{-1}$) makes little contribution to the change in the reorganization energy upon stripping the solvent, indicating that the solvent-induced electronic structure changes do not mediate the coupling of this vibration to the electronic transition.
In contrast, there is an increase in the reorganization energy due to the presence of solvent for some of the higher-frequency peaks in the spectral density (gray shaded region). To better understand why the solvent more strongly couples to certain nuclear motions, we analyzed the electron density difference upon excitation and the extent of H-bonding between the chromophore and ethanol solvent molecules.
The difference between the ground and excited-state electron density (Fig 5, top) shows that the most notable changes occur in the ring systems. However, the H-bond donating primary amine, although not exhibiting a large difference in electron density itself, provides a pathway to couple solvation effects to electron density changes in the ring system. This amine shows significant H-bonding in the molecular dynamics simulations, acting as an H-bond donor to the solvent ethanol molecules 98.5% of the time (with two H-bonds to ethanol shown in Fig 5, top), whereas the chromophore ring oxygen or nitrogen atoms act as H-bond acceptors at most 0.8% of the time and the nitrogen in the sterically hindered tertiary amine acts as an H-bond acceptor just 0.1% of the time (SI Sec. VIII). Thus, the high frequency modes that contribute to the change in the reorganization energy due to the solvent are those that involve ring modes that can receive or donate electron density via the primary amine. We note, however, that the mode at 1640 $\mathrm{cm}^{-1}$ itself does not change much in intensity upon solvation since this mode predominantly involves the amine with less contribution from the ring systems (SI Sec. III). In addition, the central ring breathing mode at 587 $\mathrm{cm}^{-1}$ (SI Sec. III), although possessing a large contribution to the change in the electron density upon excitation and hence appearing intensely in the spectral density, does not exhibit sensitivity to solvent and hence does not contribute to the change in the reorganization energy between the solvated and solvent-stripped simulations. These observations highlight the importance of considering both the locations of the changes to the electron density upon moving from the ground to the excited state and also how H-bonding sites interact with that electron density. Both of these factors affect the reorganization energy and the overall ability of solvent to control relaxation on the excited state surface.
In summary, we have shown that one can construct ML potential energy surfaces using the EqT architecture capable of generating linear and 2DES spectra for a chromophore in its condensed phase chemical environment with orders of magnitude fewer ground and excited state electronic structure calculations than would be needed to perform these calculations directly using ab initio molecular dynamics. By applying this approach to Nile blue in ethanol and comparing it to recent experiments, we show that we can obtain good agreement with both the shape of the linear, pump-probe, and multidimensional spectra by combining TDDFT electronic structure with a second-order cumulant expansion in the electronic energy gap fluctuations treatment of the spectra. By rigorously breaking down the simulated spectra into signals arising from stimulated emission and the ground-state bleach, we assessed the dynamic Stokes shift and the emergence of the shifts arising from the long-time stabilization of the electronic excited state due to nuclear motion of the chromophore and solvent. Finally, by creating an electronic energy gap surface trained on solvent-stripped electronic structure calculations, we have shown how one can break down the reorganization energy into frequency-dependent components that indicate the key regions of nuclear motions of the chromophore and solvent that couple to the electronic transition. We believe that the ability to efficiently and accurately simulate the complex interactions that contribute to multidimensional spectra enabled by these developments will lead to deeper synergies between experiment and theory going forward. 2DES has been widely applied to molecular, materials, and biological systems that rely on coupled electronic and nuclear dynamics, but specific molecular assignments to features and phenomena within these spectra has often been intractable. Through this advance, we can now simulate these measurements, enabling interpretation through the lens of ground-up theories.
Supporting Information
Additional details of the experimental methods, electronic structure calculations, training of machine learning models, molecular dynamics simulations, assessment of ML-based computational cost acceleration, normal modes that contribute to the spectral density of Nile blue in ethanol, comparison between experimental and simulated pump-probe spectra, assessment of the GSB and SE components of the simulated 2DES spectra without the laser profile, reweighting procedure to obtain the Stoke shift, and hydrogen bonding sites of Nile blue (PDF)
Data sets, EqT models, and input files to train the ground state and electronic excitation energy gap EqT models (10.5281/zenodo.15376273)
Acknowledgments
This work was funded by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences (DE-SC0020203 to C.M.I. and T.E.M.) and the National Science Foundation Grant No. CHE-2154291 (to T.E.M.). M.S., A.L., and G.S.S.-C. were funded by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Division of Chemical Sciences, Geosciences, and Biosciences under Award #DE-SC0018097 to G.S.S.-C. M.S.C. acknowledges support from the Simons Center for Computational Physical Chemistry at New York University. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231 using NERSC Awards BES-ERCAP0019987 and BES-ERCAP0023755. A.M.C. was funded by an Early Career Award in the CPIMS program in the Chemical Sciences, Geosciences, and Biosciences Division of the Office of Basic Energy Sciences of the U.S. Department of Energy under Award No. DE-SC0024154. Some of the computing for this project was performed on the Sherlock cluster. We would like to thank Stanford University and the Stanford Research Computing Center for providing computational resources and support that contributed to these research results.